OpenStudy (anonymous):

i think this question has to do with logarithms: 8-3^x=-1 does anyone know how to solve it?

OpenStudy (phi):

How about add 3^x to both sides of the equation to get 8= -1+3^x now add 1 to both sides: 8+1 = -1+1+3^x or 3^x= 9 The equation says "3 to some power equals 9" So your first thought might (should?) be, write 9 as 3 to some power. 3^x= 3^2 so it looks like x=2. Try x=2 in the original equation: 8 =? -1+ 3^2 -1+9= 8 yes. it works.